OpenStudy (anonymous):

Integrate with only calc 1 techniques: sqrt((x-1)/x^5)dx

OpenStudy (anonymous):

\[\int\limits_{}^{} \sqrt{\frac{ x-1 }{ x^{5} }}dx\]

OpenStudy (anonymous):

The only thing I could even think to try to do after all my lame sub attempts was to trig sub. But this is a problem out of a section where only u-sub has been introduced.

ganeshie8 (ganeshie8):

\(\large \int\limits_{}^{} \sqrt{\frac{ x-1 }{ x^{5} }}dx \) \(\large \int\limits_{}^{} \frac{1}{x^2}\sqrt{\frac{ x-1 }{ x }}dx \)

ganeshie8 (ganeshie8):

sub u = 1/x

OpenStudy (anonymous):

not (1- 1/x)?

ganeshie8 (ganeshie8):

thats even better :)

OpenStudy (anonymous):

Alrighty, cool. Wouldve never seen it, lol. Thanks.

hartnn (hartnn):

could you solve it entirely ? can you show us how did u solve ? if u won't mind...

hartnn (hartnn):

@Concentrationalizing @ganeshie8